# A Linear Bound on the Diameter of the Kakimizu Complex for Hyperbolic Knots

### v3

@@ 6 Annulus Surgery
  • Modify $S_{1}$ to obtain $\mathcal{S}_{1}^{\prime}$ by
    $\mathcal{S}_{1}^{\prime}:=\left(S_{1}\setminus\left((\partial A)\times I\right)\right)\cup\left(A\times\{0,1\}\right).$
  • We take the connected component of $\mathcal{S}_{1}^{\prime}$ containing the boundary $K$ and denote it by $\mathcal{S}_{1}$.
We say that $\mathcal{S}_{1}$ is obtained from $S_{1}$ by an annulus surgery along $A$ on $S_{2}$.
@@
Annulus surgery does not increase the genus, that is, $g(\mathcal{S}_{1})\leq g(S_{1})$. In particular, if $S_{1}$ has minimal genus, then $\mathcal{S}_{1}$ also has minimal genus.
@@
Observe that $\mathcal{S}^{\prime}_{1}$ is obtained from $S_{1}$ by cutting along two disjoint closed loops (creating four boundary components) and then pairwise regluing these boundary components. This process preserves the Euler characteristic:
$\chi(S_{1})=\chi(\mathcal{S}^{\prime}_{1}).$
Moreover, since $\widetilde{A}$ lies entirely on one side of $\widetilde{\mathcal{S}_{1}}$, the resulting surface $\mathcal{S}^{\prime}_{1}$ after the surgery is orientable.
  • Connected case: if $\mathcal{S}^{\prime}_{1}$ is connected, then $\mathcal{S}_{1}=\mathcal{S}^{\prime}_{1}$. Since $\chi(S_{1})=\chi(\mathcal{S}^{\prime}_{1})$ and both surfaces are bounded by $K$, it follows that $\mathcal{S}_{1}$ and $S_{1}$ have the same genus.
  • Disconnected case: if $\mathcal{S}^{\prime}_{1}$ is disconnected, then it consists of two components: the surface $\mathcal{S}_{1}$ bounded by $K$, and a closed surface $\mathcal{S}^{c}_{1}$. The closed component $\mathcal{S}^{c}_{1}$ cannot be a sphere, since it contains at least one handle in its topology. Consequently, $g(\mathcal{S}^{c}_{1})\geq 1$. Therefore, by the Euler characteristic relation,
    $\chi(S_{1})=\chi(\mathcal{S}^{\prime}_{1})=\chi(\mathcal{S}_{1})+\chi(\mathcal{S}^{c}_{1})\leq\chi(\mathcal{S}_{1}),$
    which implies that $g(S_{1})\geq g(\mathcal{S}_{1})$. In particular, if $S_{1}$ has minimal genus, then the genus of $\mathcal{S}_{1}$ cannot be smaller than that of $S_{1}$, and hence $g(S_{1})=g(\mathcal{S}_{1})$. That is, both surfaces have minimal genus.
@@
By performing annulus surgery along $A\subset S_{2}$, we transform the pair $(S_{1},S_{2})$ into $(\mathcal{S}_{1},S_{2})$. If we view $(\widetilde{S_{2}})_{j}$ with respect to $S_{1}$ before the surgery, and with respect to $\mathcal{S}_{1}$ after the surgery, then the effect of the surgery can be described as follows: the annulus $\widetilde{A}\subset(\widetilde{S_{2}})_{j}^{i}$ is pushed out of the fundamental region (see Fig.˜6).
On the other hand, if we view $(\widetilde{S_{1}})_{i}$ with respect to $S_{2}$ before the surgery, and view $(\widetilde{\mathcal{S}_{1}})_{i}$ with respect to $S_{2}$ after the surgery, then the transition from $(\widetilde{S_{1}})_{i}$ to $(\widetilde{\mathcal{S}_{1}})_{i}$ may be regarded as cutting the surface along two loops and then regluing them in such a way that the resulting surface locally lies in different fundamental regions (if the surgery produces disconnected components, we further discard the closed component) (see Fig.˜7).
@@
Let $\mathcal{S}_{1}$ be the Seifert surface obtained from $S_{1}$ by annulus surgery along $A\subset S_{2}$. Then for any $i,j\in\mathbb{Z}$, we have
$\chi\bigl(\widetilde{E}_{i}(\mathcal{S}_{1})\cap(\widetilde{S_{2}})_{j}\bigr)=\chi\bigl((\widetilde{S_{2}})_{j}^{i}\bigr)$ (6.1)
@@
$\chi\bigl(\widetilde{E}_{j}(S_{2})\cap(\widetilde{\mathcal{S}_{1}})_{i}\bigr)\geq\chi\bigl((\widetilde{S_{1}})_{i}^{j}\bigr).$ (6.2)
@@
If we view $(\widetilde{S_{2}})_{j}$ with respect to $S_{1}$ before the surgery, and with respect to $\mathcal{S}_{1}$ after the surgery, then the effect of the surgery can be described as follows: the annulus $\widetilde{A}\subset(\widetilde{S_{2}})_{j}^{i}$ is pushed out of the fundamental region. Consequently, we have $\chi\bigl(\widetilde{E}_{i}(\mathcal{S}_{1})\cap(\widetilde{S_{2}})_{j}\bigr)=\chi\bigl((\widetilde{S_{2}})_{j}^{i}\bigr)$.
On the other hand, if we view $(\widetilde{S_{1}})_{i}$ with respect to $S_{2}$ before the surgery, and view $(\widetilde{\mathcal{S}_{1}})_{i}$ with respect to $S_{2}$ after the surgery, then the transition from $(\widetilde{S_{1}})_{i}$ to $(\widetilde{\mathcal{S}_{1}})_{i}$ may be regarded as cutting the surface along two loops and then regluing them in such a way that the resulting surface locally lies in different fundamental regions. Up to this stage, the Euler characteristic is unchanged. If this operation produces disconnected components, we further discard any closed components; we obtain $\chi\bigl(\widetilde{E}_{j}(S_{2})\cap(\widetilde{\mathcal{S}_{1}})_{i}\bigr)\geq\chi\bigl((\widetilde{S_{1}})_{i}^{j}\bigr).$
Now consider the case where $S_{1}$ has minimal genus. If any closed component appears during the annulus surgery, then it must be a torus, whose Euler characteristic is zero. Otherwise, the retained component (which is a Seifert surface) would have strictly smaller genus than $S_{1}$, which contradicts the assumption that $S_{1}$ has minimal genus. Therefore, in this case, annulus surgery does not change the Euler characteristic, and we have $\chi\bigl(\widetilde{E}_{j}(S_{2})\cap(\widetilde{\mathcal{S}_{1}})_{i}\bigr)=\chi\bigl((\widetilde{S_{1}})_{i}^{j}\bigr)$.
Annulus surgery can be performed not only along a $U$-shaped annulus $A\subset S_{2}$, but also along a $U$-shaped annulus $A\subset S_{1}$. By symmetry, the conclusions of Proposition˜6.4 and Proposition˜6.5 remain valid in this case.
@@ 7 Central Zone
The main result of this chapter is to prove the following inequality concerning Euler characteristics:
@@
The reason for introducing the sum $\widehat{\chi}(h)$ is the following. By Proposition˜5.7 and Proposition˜5.10, we have
@@
Suppose that either $(\widetilde{S_{1}})_{0}^{-h-1}$ or $(\widetilde{S_{2}})_{0}^{h}$ contains at least one $U$-shaped annulus. By performing annulus surgery along such a $U$-shaped annulus, we obtain a new Seifert surface. For simplicity of notation, and since this procedure will be applied repeatedly, we continue to denote the resulting surface by $S_{1}$ or $S_{2}$, respectively.
This surgical procedure has two crucial effects. First, it preserves or decreases the genus of the surface, as established in Proposition˜6.4; more precisely, when restricted to each region, it preserves or increases the Euler characteristic, as established in Proposition˜6.5. Second, it reduces the number of boundary components of $(\widetilde{S_{1}})_{0}^{-h-1}$ and $(\widetilde{S_{2}})_{0}^{h}$ by exactly two, respectively.
After each surgery, we check again whether the new $(\widetilde{S_{1}})_{0}^{-h-1}$ or $(\widetilde{S_{2}})_{0}^{h}$ contains any $U$-shaped annuli. If so, we perform the surgery again on the corresponding surface. Since each surgery reduces the number of boundary components of $(\widetilde{S_{1}})_{0}^{-h-1}$ and $(\widetilde{S_{2}})_{0}^{h}$, after finitely many steps, both of them will contain only $I$-shaped annuli (or may be empty) and no $U$-shaped ones.
@@
    In this case, by the construction of the annulus surgeries, $(\widetilde{\mathcal{S}_{1}})_{0}$ is connected. Moreover, by Proposition˜6.4, since $S_{1}$ has minimal genus, the surface $\mathcal{S}_{1}$ also has minimal genus. By Proposition˜6.5 and the minimal-genus property of $S_{1}$, one of the following must hold: either for every integer $h_{1}^{\prime}<h$, $\chi((\widetilde{S_{1}})_{0}^{-h_{1}^{\prime}-1})=\chi((\widetilde{\mathcal{S}_{1}})_{0}^{-h_{1}^{\prime}-1})=\chi(\varnothing)=0$, or for every integer $h_{2}^{\prime}>h$, $\chi((\widetilde{S_{1}})_{0}^{-h_{2}^{\prime}-1})=\chi((\widetilde{\mathcal{S}_{1}})_{0}^{-h_{2}^{\prime}-1})=\chi(\varnothing)=0$. This completes the proof.
@@
    By appropriately ordering these components, we can ensure that for any $1\leq t\leq T$, we have $(\widetilde{\mathcal{C}}_{0,-h-1})_{t}$ and $(\widetilde{\mathcal{C}}_{0,-h})_{t}$ are $\mathbb{S}^{1}$ components connected by an annulus $\widetilde{A}_{t}((\widetilde{\mathcal{S}_{1}})_{0}^{-h-1})\subset(\widetilde{\mathcal{S}_{1}})_{0}^{-h-1}$ (see Fig.˜8). Moreover, for $h^{\prime}\in\{h,h+1\}$, we have $p((\widetilde{\mathcal{C}}_{0,-h^{\prime}})_{t})=p((\widetilde{\mathcal{C}}_{h^{\prime},0})_{t})$. In this setting, there exists a bijection $\sigma:\{1,\cdots,T\}\rightarrow\{1,\cdots,T\}$ such that for each $1\leq t\leq T$, we have $(\widetilde{\mathcal{C}}_{h,0})_{t}$ and $(\widetilde{\mathcal{C}}_{h+1,0})_{\sigma(t)}$ are connected by an annulus $\widetilde{A}_{t}((\widetilde{\mathcal{S}_{2}})_{0}^{h})\subset(\widetilde{\mathcal{S}_{2}})_{0}^{h}$.
@@
      This contradicts the incompressibility of the minimal-genus Seifert surface $\mathcal{S}_{1}$. Indeed, since $S_{1}$ has minimal genus, the surface $\mathcal{S}_{1}$ obtained from it also has minimal genus (see Proposition˜6.4).
@@
      The condition $\mathrm{ab}\circ f_{*}(\alpha)=0$ forces $\alpha$ to be isotopic to a longitude on $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$. For the case $r=1$, the torus $\mathbb{T}$ splits into two annuli $p\left(\widetilde{A}_{1}((\widetilde{\mathcal{S}_{1}})_{0}^{-h-1})\right)\subset\mathcal{S}_{1}$ and $p\left(\widetilde{A}_{1}((\widetilde{\mathcal{S}_{2}})_{0}^{h})\right)\subset\mathcal{S}_{2}$ by Eq.˜7.2. Thus by Lemma˜4.1, we conclude that $\mathbb{T}$ is not boundary-parallel.
      For the case $r>1$, we observe that $\mathrm{ab}\circ f_{*}(\pi_{1}(\mathbb{T}))=r\mathbb{Z}$. However, if $\mathbb{T}$ were boundary-parallel, this would require $\mathrm{ab}\circ f_{*}(\pi_{1}(\mathbb{T}))=\mathbb{Z}$, leading to a contradiction.
@@
In this chapter, we complete the proof of Theorem˜1.1.
@@
By Theorem˜7.1 we have $\chi\bigl((\widetilde{S_{1}})_{0}^{-h-1}\bigr)=0$ for all $h\geq h_{4}+1$. If $h_{4}+1<h_{2}$, since $(\widetilde{S_{1}})_{0}$ is connected, there exists an annulus $\widetilde{A}_{1}\subset(\widetilde{S_{1}})_{0}$ whose two boundary components lie on $(\widetilde{S_{2}})_{-(h_{4}+1)}$ and $(\widetilde{S_{2}})_{-h_{2}}$, respectively (see the upper panel of Fig.˜9). If $h_{4}+1=h_{2}$, we define $\widetilde{A}_{1}$ to be an arbitrarily chosen connected component of $(\widetilde{S_{1}})_{0}\cap(\widetilde{S_{2}})_{-h_{2}}$, viewed as a degenerate annulus consisting of a single loop. In either case, $\widetilde{A}_{1}$ intersects $(\widetilde{S_{2}})_{-h}$ in a collection of loops for each $h$ with $h_{4}+1\leq h\leq h_{2}$. For each such $h$, we choose an arbitrary loop $\widetilde{C}_{h}\subset\widetilde{A}_{1}\cap(\widetilde{S_{2}})_{-h}$, and define the collection of chosen loops
@@
Similarly, if $h_{1}+1<h_{3}$, there exists an annulus $\widetilde{A}_{2}\subset(\widetilde{S_{1}})_{0}$ whose two boundary components lie on $(\widetilde{S_{2}})_{-(h_{1}+1)}$ and $(\widetilde{S_{2}})_{-h_{3}}$, respectively (see the upper panel of Fig.˜9). If $h_{1}+1=h_{3}$, we define $\widetilde{A}_{2}$ to be an arbitrarily chosen connected component of $(\widetilde{S_{1}})_{0}\cap(\widetilde{S_{2}})_{-h_{3}}$. In either case, $\widetilde{A}_{2}$ intersects $(\widetilde{S_{2}})_{-h}$ for all $h$ with $h_{1}+1\leq h\leq h_{3}$. As before, we choose an arbitrary loop $\widetilde{C}_{h}\subset\widetilde{A}_{2}\cap(\widetilde{S_{2}})_{-h}$ for each $h$, and define
@@
We argue by contradiction. For the first statement, suppose that there exist two distinct elements $\widetilde{C}_{i},\widetilde{C}_{j}\in\widetilde{\mathcal{E}}_{1}$ such that their images $p(\widetilde{C}_{i})$ and $p(\widetilde{C}_{j})$ are isotopic on $S_{2}$. Then $p(\widetilde{C}_{i})$ and $p(\widetilde{C}_{j})$ bound annuli on both $S_{1}$ and $S_{2}$, and these two annuli together form an embedded torus. Using the same argument as in Case 2 of the proof of Theorem˜7.1, together with Lemma˜4.1, we conclude that this torus is essential. This contradicts the assumption that $K$ is atoroidal. Hence the assumption is false, and the claim 1 follows. The argument for $\widetilde{\mathcal{E}}_{2}$ is identical.
@@
Since $p(\widetilde{C}_{i})$ and $p(\widetilde{C}_{j})$ are isotopic on $S_{2}$, the loops $\tau^{i}(\widetilde{C}_{i})$ and $\tau^{j}(\widetilde{C}_{j})$ bound an annulus in $(\widetilde{S}_{2})_{0}$, which we denote by $\widetilde{A}_{3}$ (see the bottom panel of Fig.˜9). Moreover, $\widetilde{A}_{3}$ intersects $(\widetilde{S}_{1})_{h}$ for all $i\leq h\leq j$, and these intersections consist of ji+1j-i+1 loop components.
@@
On the other hand, Sakuma–Shackleton showed that when $g=1$, the diameter of the Kakimizu complex satisfies $\operatorname{diam}(MS(K))\leq 2$ [SS09, Corollary 1.3]. Combining these results, we obtain the following corollary.

### v4

@@ 6 Annulus Surgery
  • Modify $S_{1}$ to obtain $S_{1}^{\prime}$ by
    $S_{1}^{\prime}:=\left(S_{1}\setminus\left((\partial A)\times I\right)\right)\cup\left(A\times\{0,1\}\right).$
  • We take the connected component of $S_{1}^{\prime}$ containing the boundary $K$ and denote it by $s(S_{1})$.
We say that $s(S_{1})$ is obtained from $S_{1}$ by an annulus surgery along $A$ on $S_{2}$.
@@
This figure also explains why annulus surgery can only be performed on $U$-shaped annuli, but not on $I$-shaped annuli. Since $\widetilde{A}$ lies entirely on one side of $\widetilde{S_{1}}$, the resulting surface after the surgery is orientable. If one were to perform the surgery on an $I$-shaped annulus instead, the resulting surface would be non-orientable.
Remark 6.4.
The definition of annulus surgery is not restricted to being performed along a $U$-shaped annulus $A\subset S_{2}$, but can also be extended to a $U$-shaped annulus $A\subset S_{1}$. For convenience of notation, we view annulus surgery as a transformation from a pair of surfaces to another pair of surfaces:
$s(S_{1},S_{2})=(s^{A}(S_{1}),s^{A}(S_{2})).$
Here, the superscript $A$ indicates that the surgery is performed along the annulus $A$, which may lie in either $S_{1}$ or $S_{2}$. If $A\subset S_{2}$, we define $s^{A}(S_{2}):=S_{2}$, whereas if $A\subset S_{1}$, we define $s^{A}(S_{1}):=S_{1}$. When no confusion arises, we will sometimes omit the superscript.
Moreover, we denote
$\widetilde{(s^{A}(S_{1}))}_{i}^{j}:=\widetilde{E}_{j}(s^{A}(S_{2}))\cap(\widetilde{s^{A}(S_{1})})_{i}\quad\text{and}\quad\widetilde{(s^{A}(S_{2}))}_{j}^{i}:=\widetilde{E}_{i}(s^{A}(S_{1}))\cap(\widetilde{s^{A}(S_{2})})_{j}.$
@@
Let $A$ be an annulus on $S_{2}$ such that its lift $\widetilde{A}$ is a $U$-shaped annulus. Then,
$g(s(S_{1}))\leq g(S_{1}).$
In particular, if $S_{1}$ has minimal genus, then $g(s(S_{1}))=g(S_{1})$; that is, $s(S_{1})$ also has minimal genus.
@@
Observe that $S^{\prime}_{1}$ is obtained from $S_{1}$ by cutting along two disjoint closed loops (creating four boundary components) and then pairwise regluing these boundary components. This process preserves the Euler characteristic:
$\chi(S_{1})=\chi(S^{\prime}_{1}).$
Moreover, since $\widetilde{A}$ lies entirely on one side of $\widetilde{S_{1}}$, the resulting surface $S^{\prime}_{1}$ after the surgery is orientable.
  • Connected case: if $S^{\prime}_{1}$ is connected, then $s(S_{1})=S^{\prime}_{1}$. Since $\chi(S_{1})=\chi(S^{\prime}_{1})$ and both surfaces have only one boundary component, namely $K$, it follows that $s(S_{1})$ and $S_{1}$ have the same genus.
  • Disconnected case: if $S^{\prime}_{1}$ is disconnected, then it consists of two components: the surface $s(S_{1})$ bounded by $K$, and a closed surface $S^{{}^{\prime\prime}}_{1}:=S^{\prime}_{1}\setminus s(S_{1})$. The closed component $S^{{}^{\prime\prime}}_{1}$ cannot be a sphere, since it contains at least one handle in its topology. Consequently, $g(S^{{}^{\prime\prime}}_{1})\geq 1$. Therefore, by the Euler characteristic relation,
    $\chi(S_{1})=\chi(S^{\prime}_{1})=\chi(s(S_{1}))+\chi(S^{{}^{\prime\prime}}_{1})\leq\chi(s(S_{1})),$
    which implies that $g(S_{1})\geq g(s(S_{1}))$. In particular, if $S_{1}$ has minimal genus, then the genus of $s(S_{1})$ cannot be smaller than that of $S_{1}$, and hence $g(S_{1})=g(s(S_{1}))$. That is, both surfaces have minimal genus.
@@
By performing annulus surgery along $A\subset S_{2}$, we transform the pair $(S_{1},S_{2})$ into $s(S_{1},S_{2})$. If we view $(\widetilde{S_{2}})_{j}$ with respect to $\{(\widetilde{S_{1}})_{i}\}_{i\in\mathbb{Z}}$ before the surgery, and view $(\widetilde{s(S_{2})})_{j}=(\widetilde{S_{2}})_{j}$ with respect to $\{(\widetilde{s(S_{1})})_{i}\}_{i\in\mathbb{Z}}$ after the surgery, then the effect of the surgery can be described as follows: the annulus $\widetilde{A}$ is pushed out of the fundamental region (see Fig.˜6).
On the other hand, if we view $(\widetilde{S_{1}})_{i}$ with respect to $\{(\widetilde{S_{2}})_{j}\}_{j\in\mathbb{Z}}$ before the surgery, and view $(\widetilde{s(S_{1})})_{i}$ with respect to $\{(\widetilde{s(S_{2})})_{j}\}_{j\in\mathbb{Z}}=\{(\widetilde{S_{2}})_{j}\}_{j\in\mathbb{Z}}$ after the surgery, then the surgery may be regarded as cutting the surface $(\widetilde{S_{1}})_{i}$ along two loops and then regluing them in such a way that the resulting surface locally lies in different fundamental regions (if the surgery produces disconnected components, we further discard the closed component) (see Fig.˜7).
@@
Let $(s(S_{1}),s(S_{2}))$ be the pair of Seifert surfaces obtained from $(S_{1},S_{2})$ by annulus surgery along $A\subset S_{2}$. Recall that $\widetilde{(s(S_{1}))}_{i}^{j}:=\widetilde{E}_{j}(s(S_{2}))\cap(\widetilde{s(S_{1})})_{i}$ and $\widetilde{(s(S_{2}))}_{j}^{i}:=\widetilde{E}_{i}(s(S_{1}))\cap(\widetilde{s(S_{2})})_{j}$. Then for any $i,j\in\mathbb{Z}$, we have
$\chi\bigl((\widetilde{s(S_{2})})_{j}^{i}\bigr)=\chi\bigl((\widetilde{S_{2}})_{j}^{i}\bigr)$ (6.1)
@@
$\chi\bigl((\widetilde{s(S_{1})})_{i}^{j}\bigr)\geq\chi\bigl((\widetilde{S_{1}})_{i}^{j}\bigr).$ (6.2)
@@
If we view $(\widetilde{S_{2}})_{j}$ with respect to $\{(\widetilde{S_{1}})_{i}\}_{i\in\mathbb{Z}}$ before the surgery, and view $(\widetilde{s(S_{2})})_{j}$ with respect to $\{(\widetilde{s(S_{1})})_{i}\}_{i\in\mathbb{Z}}$ after the surgery, then the effect of the surgery can be described as follows: the annulus $\widetilde{A}\subset(\widetilde{S_{2}})_{j}^{i}$ is pushed out of the fundamental region. Consequently, we have $\chi\bigl((\widetilde{s(S_{2})})_{j}^{i}\bigr)=\chi\bigl((\widetilde{S_{2}})_{j}^{i}\bigr)$.
On the other hand, if we view $(\widetilde{S_{1}})_{i}$ with respect to $\{(\widetilde{S_{2}})_{j}\}_{j\in\mathbb{Z}}$ before the surgery, and view $(\widetilde{s(S_{1})})_{i}$ with respect to $\{(\widetilde{s(S_{2})})_{j}\}_{j\in\mathbb{Z}}=\{(\widetilde{S_{2}})_{j}\}_{j\in\mathbb{Z}}$ after the surgery, then the surgery may be regarded as cutting the surface $(\widetilde{S_{1}})_{i}$ along two loops and then regluing them in such a way that the resulting surface locally lies in different fundamental regions. Up to this stage, the Euler characteristic is unchanged. If this operation produces disconnected components, we further discard any closed components; we obtain $\chi\bigl((\widetilde{s(S_{1})})_{i}^{j}\bigr)\geq\chi\bigl((\widetilde{S_{1}})_{i}^{j}\bigr).$
Now consider the case where $S_{1}$ has minimal genus. If any closed component appears during the annulus surgery, then it must be a torus, whose Euler characteristic is zero. Otherwise, the retained component (which is a Seifert surface) would have strictly smaller genus than $S_{1}$, which contradicts the assumption that $S_{1}$ has minimal genus. Therefore, in this case, annulus surgery does not change the Euler characteristic, and we have $\chi\bigl((\widetilde{s(S_{1})})_{i}^{j}\bigr)=\chi\bigl((\widetilde{S_{1}})_{i}^{j}\bigr)$.
Annulus surgery can be performed not only along a $U$-shaped annulus $A\subset S_{2}$, but also along a $U$-shaped annulus $A\subset S_{1}$. By symmetry, the conclusions of Proposition˜6.5 and Proposition˜6.6 remain valid in this case.
@@ 7 Central Zone
The main result of this section is to prove the following inequality concerning Euler characteristics:
@@
The reason for introducing the sum $\widehat{\chi}(h)$ is the following. By Proposition˜5.8 and Proposition˜5.10, we have
@@
Suppose that either $(\widetilde{S_{1}})_{0}^{-h-1}$ or $(\widetilde{S_{2}})_{0}^{h}$ contains at least one $U$-shaped annulus. By performing annulus surgery along such a $U$-shaped annulus $A_{1}$, we obtain $(\widetilde{s^{A_{1}}(S_{1})})_{0}^{-h-1}$ and $(\widetilde{s^{A_{1}}(S_{2})})_{0}^{h}$.
This surgical procedure has two crucial effects. First, it preserves or decreases the genus of the surface, as established in Proposition˜6.5; more precisely, when restricted to each region, it preserves or increases the Euler characteristic, as established in Proposition˜6.6. Second, it reduces the number of boundary components of $(\widetilde{S_{1}})_{0}^{-h-1}$ and $(\widetilde{S_{2}})_{0}^{h}$ by exactly two, respectively.
Since annulus surgery preserves or increases the Euler characteristic, and the Euler characteristics of $(\widetilde{s^{A_{1}}(S_{1})})_{0}^{-h-1}$ and $(\widetilde{s^{A_{1}}(S_{2})})_{0}^{h}$ cannot increase any further, they must still be equal to 0. Equivalently, each of them remains a disjoint union of annuli. We now check whether $(\widetilde{s^{A_{1}}(S_{1})})_{0}^{-h-1}$ and $(\widetilde{s^{A_{1}}(S_{2})})_{0}^{h}$ contain any $U$-shaped annuli. If so, we perform the surgery again on the corresponding surface, we obtain $(\widetilde{s^{A_{2}}s^{A_{1}}(S_{1})})_{0}^{-h-1}$ and $(\widetilde{s^{A_{2}}s^{A_{1}}(S_{2})})_{0}^{h}$, and so on. Since each surgery reduces the number of boundary components of $(\widetilde{S_{1}})_{0}^{-h-1}$ and $(\widetilde{S_{2}})_{0}^{h}$, after finitely many steps, both of $(\widetilde{s^{A_{m}}\cdots s^{A_{2}}s^{A_{1}}(S_{1})})_{0}^{-h-1}$ and $(\widetilde{s^{A_{m}}\cdots s^{A_{2}}s^{A_{1}}(S_{2})})_{0}^{h}$ will contain only $I$-shaped annuli (or may be empty) and no $U$-shaped ones (see Fig.˜8 for an example).
Figure 8: Performing multiple annulus surgeries on $(\widetilde{S_{1}})_{0}^{-h-1}$ and $(\widetilde{S_{2}})_{0}^{h}$.
@@
    In this case, by the construction of the annulus surgeries, $(\widetilde{\mathcal{S}_{1}})_{0}$ is connected. Moreover, by Proposition˜6.5, since $S_{1}$ has minimal genus, the surface $\mathcal{S}_{1}$ also has minimal genus. By Proposition˜6.6 and the minimal-genus property of $S_{1}$, one of the following must hold: either for every integer $h_{1}^{\prime}<h$, $\chi((\widetilde{S_{1}})_{0}^{-h_{1}^{\prime}-1})=\chi((\widetilde{\mathcal{S}_{1}})_{0}^{-h_{1}^{\prime}-1})=\chi(\varnothing)=0$, or for every integer $h_{2}^{\prime}>h$, $\chi((\widetilde{S_{1}})_{0}^{-h_{2}^{\prime}-1})=\chi((\widetilde{\mathcal{S}_{1}})_{0}^{-h_{2}^{\prime}-1})=\chi(\varnothing)=0$. This completes the proof.
@@
    By appropriately ordering these components, we can ensure that for any $1\leq t\leq T$, we have $(\widetilde{\mathcal{C}}_{0,-h-1})_{t}$ and $(\widetilde{\mathcal{C}}_{0,-h})_{t}$ are $\mathbb{S}^{1}$ components connected by an annulus $\widetilde{A}_{t}((\widetilde{\mathcal{S}_{1}})_{0}^{-h-1})\subset(\widetilde{\mathcal{S}_{1}})_{0}^{-h-1}$ (see Fig.˜9). Moreover, for $h^{\prime}\in\{h,h+1\}$, we have $p((\widetilde{\mathcal{C}}_{0,-h^{\prime}})_{t})=p((\widetilde{\mathcal{C}}_{h^{\prime},0})_{t})$. In this setting, there exists a bijection $\sigma:\{1,\cdots,T\}\rightarrow\{1,\cdots,T\}$ such that for each $1\leq t\leq T$, we have $(\widetilde{\mathcal{C}}_{h,0})_{t}$ and $(\widetilde{\mathcal{C}}_{h+1,0})_{\sigma(t)}$ are connected by an annulus $\widetilde{A}_{t}((\widetilde{\mathcal{S}_{2}})_{0}^{h})\subset(\widetilde{\mathcal{S}_{2}})_{0}^{h}$.
@@
      This contradicts the incompressibility of the minimal-genus Seifert surface $\mathcal{S}_{1}$. Indeed, since $S_{1}$ has minimal genus, the surface $\mathcal{S}_{1}$ obtained from it also has minimal genus (see Proposition˜6.5).
@@
      The condition $\mathrm{ab}\circ\iota_{*}(\alpha)=0$ forces $\alpha$ to be isotopic to a longitude on $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$. For the case $r=1$, the torus $\mathbb{T}$ splits into two annuli $p\left(\widetilde{A}_{1}((\widetilde{\mathcal{S}_{1}})_{0}^{-h-1})\right)\subset\mathcal{S}_{1}$ and $p\left(\widetilde{A}_{1}((\widetilde{\mathcal{S}_{2}})_{0}^{h})\right)\subset\mathcal{S}_{2}$ by Eq.˜7.2. Thus by Lemma˜4.1, we conclude that $\mathbb{T}$ is not boundary-parallel.
      For the case $r>1$, we observe that $\mathrm{ab}\circ\iota_{*}(\pi_{1}(\mathbb{T}))=r\mathbb{Z}$. However, if $\mathbb{T}$ were boundary-parallel, this would require $\mathrm{ab}\circ\iota_{*}(\pi_{1}(\mathbb{T}))=\mathbb{Z}$, leading to a contradiction.
@@
In this section, we complete the proof of Theorem˜1.1.
@@
By Theorem˜7.1 we have $\chi\bigl((\widetilde{S_{1}})_{0}^{-h-1}\bigr)=0$ for all $h\geq h_{4}+1$. If $h_{4}+1<h_{2}$, since $(\widetilde{S_{1}})_{0}$ is connected, there exists an annulus $\widetilde{A}_{1}\subset(\widetilde{S_{1}})_{0}$ whose two boundary components lie on $(\widetilde{S_{2}})_{-(h_{4}+1)}$ and $(\widetilde{S_{2}})_{-h_{2}}$, respectively (see the upper panel of Fig.˜10). If $h_{4}+1=h_{2}$, we define $\widetilde{A}_{1}$ to be an arbitrarily chosen connected component of $(\widetilde{S_{1}})_{0}\cap(\widetilde{S_{2}})_{-h_{2}}$, viewed as a degenerate annulus consisting of a single loop. In either case, $\widetilde{A}_{1}$ intersects $(\widetilde{S_{2}})_{-h}$ in a collection of loops for each $h$ with $h_{4}+1\leq h\leq h_{2}$. For each such $h$, we choose an arbitrary loop $\widetilde{C}_{h}\subset\widetilde{A}_{1}\cap(\widetilde{S_{2}})_{-h}$, and define the collection of chosen loops
@@
Similarly, if $h_{1}+1<h_{3}$, there exists an annulus $\widetilde{A}_{2}\subset(\widetilde{S_{1}})_{0}$ whose two boundary components lie on $(\widetilde{S_{2}})_{-(h_{1}+1)}$ and $(\widetilde{S_{2}})_{-h_{3}}$, respectively (see the upper panel of Fig.˜10). If $h_{1}+1=h_{3}$, we define $\widetilde{A}_{2}$ to be an arbitrarily chosen connected component of $(\widetilde{S_{1}})_{0}\cap(\widetilde{S_{2}})_{-h_{3}}$. In either case, $\widetilde{A}_{2}$ intersects $(\widetilde{S_{2}})_{-h}$ for all $h$ with $h_{1}+1\leq h\leq h_{3}$. As before, we choose an arbitrary loop $\widetilde{C}_{h}\subset\widetilde{A}_{2}\cap(\widetilde{S_{2}})_{-h}$ for each $h$, and define
@@
We argue by contradiction. For the first statement, suppose that there exist two distinct elements $\widetilde{C}_{i},\widetilde{C}_{j}\in\widetilde{\mathcal{E}}_{1}$ such that their images $p(\widetilde{C}_{i})$ and $p(\widetilde{C}_{j})$ are isotopic on $S_{2}$. Then $p(\widetilde{C}_{i})$ and $p(\widetilde{C}_{j})$ bound annuli on both $S_{1}$ and $S_{2}$, and these two annuli together form an embedded torus. Using the same argument as in Case 2 of the proof of Theorem˜7.1, together with Lemma˜4.1, we conclude that this torus is essential. This contradicts the assumption that $K$ is atoroidal. Hence the assumption is false, and claim 1 follows. The argument for $\widetilde{\mathcal{E}}_{2}$ is identical.
@@
Since $p(\widetilde{C}_{i})$ and $p(\widetilde{C}_{j})$ are isotopic on $S_{2}$, the loops $\tau^{i}(\widetilde{C}_{i})$ and $\tau^{j}(\widetilde{C}_{j})$ bound an annulus in $(\widetilde{S}_{2})_{0}$, which we denote by $\widetilde{A}_{3}$ (see the bottom panel of Fig.˜10). Moreover, $\widetilde{A}_{3}$ intersects $(\widetilde{S}_{1})_{h}$ for all $i\leq h\leq j$, and these intersections consist of ji+1j-i+1 loop components.
@@
On the other hand, Sakuma–Shackleton showed that when $g=1$, the diameter of the Kakimizu complex satisfies $\operatorname{diam}(MS(K))\leq 2$ [11, Corollary 1.3]. Combining these results, we obtain the following corollary.